OpenStudy (anonymous):

A dance CD has 12 songs on it - 9 are slow, and 3 are fast. When the DJ at a dance plays a song it is not played again. All songs from the CD are played at random. What is the probability that the first two songs played are slow songs? any one know this?

OpenStudy (aum):

Out of 12 total songs, 9 are slow. If a song is picked at random, what is the probability the first song played is slow?

OpenStudy (anonymous):

can you explain? idk what to do

OpenStudy (aum):

Out of 12 total songs, 9 are slow. If a song is picked at random, the probability that it is a slow song is 9/12. Now there are 11 total songs left (because songs are not repeated) out of which 2 are slow. Now if a song is picked at random, the probability that it is a slow song is 8/11. So the probability the first song is slow AND the second song is also slow is: 9/12 * 8/11 = 72 / (12*11) = 6/11